# Deep Neural Network-Based Voltage Prediction for Alkali-Metal-Ion Battery Materials

### v1

@@ 3 Results and discussion
The loss during the training of the model was evaluated as shown in Figure 3(a). At the beginning of training (epochs close to 0), both training and testing losses are high, which is expected because the model starts with randomly initialized weights. It has not yet learned to map the input features to the target output effectively. In the first few epochs, both the training and testing losses drop steeply. This indicates that the model is learning rapidly and adjusting its weights effectively to minimize the error. As training progresses, the loss values for both training and testing data begin to plateau. This suggests that the model has learnt the underlying patterns in the data and is approaching convergence. The training loss and testing loss closely follow each other throughout the training process. This is a strong indicator that the model is generalizing well to unseen data and is not overfitting. At the final epochs (e.g., around 1000), both training and testing losses remain low and stable, further confirming that the model has achieved a good fit without overfitting or underfitting. The convergence of training and training losses demonstrates the model’s capability to learn and generalize effectively. The lack of significant divergence between the two curves suggests that the chosen hyperparameters (learning rate, architecture, etc.) are appropriate. The alignment of the testing loss with the training loss highlights that the model is not overly complex and has a good balance of bias and variance.
@@
The R2 and MSE value of training and testing presented in Figure 3(b), shows a near-perfect R2 and very low MSE for training set indicates that the model has successfully learned the patterns in the training data. The test R2 value of 0.906 shows that the model captures most of the variance in the test data. Some scatter points deviate significantly from the ideal fit line, especially at higher actual voltage values. These deviations indicate instances where the model struggle to predict accurately. The higher MSE on the test set (0.182 vs. 0.009, Figure 3(d) reflects that very small prediction errors could be present on unseen data, which is common in machine learning models. The model demonstrates strong learning on the training set and reasonable generalization on the test set, as evidenced by the high test R2 and close alignment of many points with the ideal fit line ($y=x$). Minor overfitting may exist, but the model can still make accurate predictions for the majority of the test data. From the MSE and MAE values (0.182 and 0.24V respectively), it can be concluded that our DNN model shows excellent performance as compared to the earlier reported ML models, with the respective MAE 32, 31, 33 shown in Table 1. The error bar plot in Figure 3(d) compares the mean error metrics of our model. It shows a smaller MAE with respect to the previously reported MAE values by Joshi et al.31 and Moses et al.32
The residuals appear fairly symmetric around the zero line (Figure 3(c)), which suggests that the model does not have significant bias toward overestimating or underestimating the actual values. Most residuals are clustered close to the zero line, indicating that the model’s predictions are generally accurate.
@@
DNNs 0.81 0.43 Joshi et al.31
DNNs 0.83 0.39 Moses et al.32
GNNs - - - 0.34 Louis et al.33
@@
3.1 Validation of the model
Model validation is crucial for testing its excellent performance verification and evaluation. We validate our DNN model by considering experimentally observed and commercially available battery electrode material for both Li and Na-ion battery (see Table 2). We obtained a R2 and MSE values of 0.991 and 0.012V respectively, as presented in Figure 4 which suggests the excellency of our model.
@@ 3.1 Validation of the model
The list of materials used for the validation are shown in Table 2 along with voltage data obtained from DFT, our DNN model, and experiment. The absolute deviation of our model predicted voltage with experimental and DFT voltages are evaluated through the following expressions: $\Delta V_{1}=|V_{DFT}-V_{DNN}|$ and $\Delta V_{2}=|V_{expt.}-V_{DNN}|$ respectively, and tabulated in Table 2.
@@
LiCoO2 4.155 3.808 3.803 0.005 0.297
LiFePO4 3.556 3.847 3.845 0.002 0.345
LiMnPO4 4.157 4.578 4.207 0.371 0.107
LiCoPO4 4.858 4.770 4.753 0.017 0.047
LiNiO2 3.8556 4.099 4.079 0.020 0.229
LiMn2O4 4.1559 4.577 4.545 0.032 0.395
LiTiS2 2.160 1.875 1.866 0.009 0.234
LiNiPO4 5.361 5.983 5.957 0.026 0.657
Li3V2(PO4)3 3.862 3.708 3.716 0.008 0.084
Na3V2(PO4)3 3.462 3.349 3.361 0.012 0.039
NaCoO2 2.863, 64 3.056 3.043 0.013 0.243
NaNiO2 3.065 3.315 3.277 0.038 0.277
NaTiO2 1.566 1.106 1.104 0.002 0.396
NaFePO4 3.067 3.072 3.046 0.026 0.046
@@
Our model is developed based on DFT data of battery materials and its associated features, ensuring that the deviation ($\Delta V_{1}$) between our DNN model’s predicted voltage and the DFT-calculated voltage remains minimal for nearly all validation materials. In contrast, a larger deviation ($\Delta V_{2}$) is observed when comparing the predicted values to experimental data. This discrepancy is expected, as the voltage of battery materials is influenced by various experimental conditions such as humidity, temperature, and pressure—factors that are not accounted for in the DFT dataset. Despite these challenges, our model demonstrates the ability to predict experimental voltages with minimal error, as evidenced by several cases presented in Table 2.
3.2 The Rise of Na-Ion Batteries in Energy Storage
Researchers are increasingly focusing on sodium-ion (Na-ion) batteries as a promising alternative to lithium-ion (Li-ion) technology due to several key advantages in cost, resource availability, and sustainability. One of the most significant benefits of Na-ion batteries is their lower material cost, with a theoretical cost of \$40-77/kWh, making them 30% cheaper than LiFePO4 based lithium-ion batteries68, 69. This cost advantage is particularly crucial given the rising costs and scarcity of lithium, which has led to an increase in production expenses for Li-ion batteries, reaching an average of \$137/kWh in 2020. In terms of energy density, Na-ion batteries offer up to 290 Wh/kg for active materials and 250-375 Wh/L volumetric energy density, making them competitive for energy storage applications. However, current practical Na-ion prototypes have a lower energy density (75-200 Wh/kg) compared to lithium-ion equivalents, but continuous research and development are expected to improve these values.
Another critical factor driving Na-ion battery research is cycle life. While early prototypes have shown faster degradation, some advanced Na-ion battery designs have demonstrated 4,000 cycles, rivaling lithium-ion batteries68, 70. This long cycle life is particularly attractive for applications such as grid storage, where energy density is less critical than longevity and cost-effectiveness. Overall, the growing interest in Na-ion batteries stems from their potential to provide a sustainable, cost-effective, and scalable energy storage solution without the resource limitations associated with lithium. As research progresses, improvements in energy density, stability, and charge-discharge performance could make Na-ion batteries a viable alternative to Li-ion batteries for a wide range of applications, including renewable energy storage, electric vehicles, and large-scale power grids. As a result the development of Na-ion battery materials has gained significant attention in both academia and industry.
Among various Na-ion cathode materials, many industries (such as HiNa, CATL, Faradian, Tiamat, etc) has focused on layered transition metal oxides of O3 and P2-types as it emerged as highly promising due to their superior capacity compared to poly-anionic and Prussian blue-based cathodes71. However, the large ionic size of Na+ slows down the kinetics and induces complex phase transitions during insertion and extraction, leading to challenges such as poor cycle stability, low energy density, and limited rate capability. To address these challenges, researchers have explored various strategies for optimizing Na-ion cathodes. In this work, we leverage a DNN model developed using the PyTorch library to design novel O3- and P2-type layered transition metal oxide compositions.
3.3 Design Novel Cathode for Na-ion Batteries
@@
The design of new Na-ion battery cathode materials was carried out by leveraging the template crystal structures of transition metal layered oxides with the general formula NaMO2, where M represents transition or post-transition metals. These layered oxides serve as a fundamental structural framework due to their well-established electrochemical properties and stability in Na-ion battery applications. For this study, we focused on two widely studied structural types of layered transition metal oxides: O3-type and P2-type materials (see Figure 5). The O3-type materials exhibit trigonal symmetry with the $R-3m$ space group, while the P2-type materials possess hexagonal symmetry with the $P6_{3}/mmc$ space group. These two structural configurations were carefully considered in designing new compositions due to their distinct sodium diffusion pathways and electrochemical characteristics.
@@
To generate novel compositions, we explored a wide range of 3d transition metals (Cr, Ti, V, Mn, Fe, Ni, Co, Cu, Zn) along with other selected metal species (Al, Mg, Sr), aiming to achieve an optimal balance of critical electrochemical properties. The primary selection criteria for these elements were: phase stability, high voltage, long cycle life, efficient Na-ion diffusion, and fast charging capability. The Supercell package 72 was utilized to systematically explore and enumerate various element combinations within the NaMO2 framework. This computational approach enabled the creation of an extensive library of potential compositions by generating numerous atomic substitutions within the host structure. The resulting dataset includes a large number of unique chemical compositions for both O3- and P2-type cathodes, each exhibiting distinct electrochemical and structural properties.
3.4 First Principles Calculation
All density functional theory (DFT) calculations were performed using the Vienna Ab initio Simulation Package (VASP)73, 74 with spin-polarized ferromagnetic ordering. The formation energy calculations employed the generalized gradient approximation (GGA) with the Perdew–Burke–Ernzerhof (PBE) exchange–correlation functional75. For voltage calculations, we applied the Hubbard correction parameters (U) to account for self-interaction errors, and the DFT-D3 method76 was used to incorporate dispersion effects for all compounds of interest. The U parameters used for Fe, Mn, Ni, Co, Cr, and Cu were 3.9, 4.0, 6.2, 3.32, 3.7, and 5.5 eV, respectively, as obtained from Ref 77, 78. A kinetic energy cutoff of 500 eV and a $\Gamma$-centered k-point mesh of 11 × 7 × 3 for O3 and 3 x 3 x 2 for P2-type materials was found to be sufficient for achieving convergence. During structural optimization, atomic positions were allowed to relax until the Hellmann–Feynman forces were reduced below 0.01 eV/Å. The electronic minimization was performed with a convergence criterion of $10^{-6}$ eV. The new composition of charge balance structure for both O3 and P2-type was build using Supercell package 72. To generate the new compositions, here we have consider 2 x 3 x 1 supercell lattice for O3-type and 3 x 3 x 1 supercell lattice for P2-type materials.
The formation energy per formula unit (eV/f.u.) has been calculated using the below formula 79, 80:
Where $a+b+c=1$ and $M_{1}^{a}$ ,$M_{2}^{b}$ and $M_{3}^{c}$, represents the three different stoichiometry of metallic cations and their corresponding chemical potentials $\mu_{M_{1}}$, $\mu_{M_{2}}$, and $\mu_{M_{3}}$ repectively. Again $\mu_{Na}$ and $\mu_{O_{2}}$ represents the chemical potential of Na bulk and O2 molecule.
The voltage (V) of the newly designed compositions was calculated using the general formula81 for multi-component layered metal oxide which is
$V=-\frac{E_{Na_{x_{2}}MO_{2}}-E_{Na_{x_{1}}MO_{2}}-(x_{2}-x_{1})E_{Na}}{x_{2}-x_{1}}$ (4)
Where M is the multi-cations of different stoichiometry of $M_{1}^{a}$ ,$M_{2}^{b}$ and $M_{3}^{c}$; $E_{Na}$ is total energy of the Na bulk and $x_{1}$ and $x_{2}$ are the different sodium (Na) concentration during sodiation and desodiation.
A comprehensive list of the newly designed O3 and P2-type cathode materials, along with their predicted voltages (V) from our DNN model, is presented in Figure 6. The voltage predictions were obtained using our PyTorch-based DNN model, which has demonstrated superior accuracy in predicting the electrochemical performance of Na-ion cathodes. To further validate the stability and feasibility of these newly designed compositions, we performed first-principles DFT calculations. Table 3 and 4 represent the formation energy newly design material compositions, assessing their thermodynamic stability and likelihood of successful synthesis. A detailed discussion of the DFT results, including computed formation energies and structural analyses, is provided in the Supporting Information (SI) section.
@@ 3.4 First Principles Calculation
3.5 Model Predicted vs DFT Calculated Voltage
Using DFT calculations within the GGA+U framework, we have validated the voltage predictions generated by our DNN model for specific compositions: O3-Na0.67Mn0.61Ni0.28Cr0.11O2 and P2-Na0.67Mn0.5Fe0.28Cu0.22O2. These compositions, highlighted in red in Figure 6(a, b), were selected as representative validation points to avoid the computational expense of performing DFT calculations for all possible compositions. The DNN-predicted voltages for O3-Na0.67Mn0.61Ni0.28Cr0.11O2 and P2-Na0.67Mn0.5Fe0.28Cu0.22O2 are 3.79 and 3.76V, respectively, while the corresponding DFT-calculated voltages are 3.84 and 3.68V (see SI), as shown in Figure 6(a, b). The minimal errors of just 1.3% and 2.2% demonstrate an excellent agreement between the two methods. This strong correlation highlights the robustness and reliability of our machine learning model in accurately predicting voltages while significantly reducing the computational cost associated with traditional first-principles calculations. The consistency between DFT and DNN results underscore the potential of our model as a powerful tool for screening and predicting the electrochemical properties of novel sodium-ion battery materials.

### v3

@@ 4.1 DNN performance and cross-validation
The MSE loss during the model training was evaluated as shown in Figure 4(a). At the beginning of training (epochs close to 0), both training and testing losses are high, which is expected because the model starts with randomly initialized weights. It has not yet learned to map the input features to the target output effectively. In the first few epochs, both the training and the testing losses drop steeply. This indicates that the model is learning rapidly and effectively by adjusting its weights to minimize error. As training progresses, the loss values for both training and testing data begin to plateau. This suggests that the model has learned the underlying patterns in the data and is approaching convergence. Training loss and testing loss closely follow each other throughout the training process. This is a strong indicator that the model is generalizing well to unseen data and is not overfitting. At the final epochs (e.g., around 1000), both training and testing losses remain low and stable, further confirming that the model has achieved a good fit without over-fitting or under-fitting. The convergence of training and training losses demonstrates the model’s capability to learn and generalize effectively. The lack of significant divergence between the two curves suggests that the chosen hyperparameters (learning rate, architecture, etc.) are appropriate. The alignment of the test loss with the training loss highlights that the model is not overly complex and has a good balance of bias and variance. Here, the R2 score and error metrics (MSE or MAE) were initially evaluated using a single train-test split. However, to ensure superior generalization and model reliability, we have also performed a 10-fold cross-validation, as discussed below.
@@
The R2 and MSE values for training and testing presented in Figure 4 (b), show a near perfect R2 and a very low MSE for the training set, which indicate that the model has successfully learned the patterns in the training data. The test R2 value of 0.906 shows that the model captures most of the variance in the test data. Some scatter points deviate significantly from the ideal-fit line, especially at higher actual voltage values. These deviations indicate instances where the model struggles to predict accurately. The higher MSE in the test set (0.182 vs 0.009, Figure 4 (d)) reflects that very small prediction errors could be present in unseen data, which is common in machine learning models. The model demonstrates strong learning in the training set and reasonable generalization in the test set, as evidenced by the high test R2 and close alignment of many points with the ideal fit line ($y=x$). Minor overfitting may occur, but the model can still make accurate predictions for most test data. From the MSE and MAE values (0.182 and 0.24V, respectively), it can be concluded that our DNN model shows excellent performance compared to previously reported ML models, with the respective MAE 30, 29, 31 shown in Table 1. The error bar graph in Figure 4(d) compares the mean error metrics of our model. It shows a smaller MAE with respect to the MAE values previously reported by Joshi et al.29 and Moses et al.30. The residuals appear fairly symmetric around the zero line (Figure 4 (c)), suggesting that the model does not have a significant bias towards overestimating or underestimating the actual values. Most residuals are clustered close to the zero line, indicating that the model’s predictions are generally accurate.
@@
DNNs 0.81 0.43 Joshi et al.29
DNNs 0.83 0.39 Moses et al.30
GNNs - - - 0.34 Louis et al.31
@@
In addition to the DNN model, we trained several traditional regression-based machine learning models, including random forest regression (RFR), support vector regression (SVR), and gradient booster regression (GBR), to predict the average voltage of battery materials. Figure 5 illustrates the performance of these models in the training and testing phases, with the corresponding coefficient of determination R2 and the MSE values. Among the models evaluated, RFR achieved the highest predictive accuracy on the test dataset, yielding a R2 value of 0.813. Despite the relatively strong performance of the RFR model compared to SVR and GBR, it was outperformed by the DNN model, as shown in 4(b). The DNN model, implemented using the PyTorch framework, demonstrated a superior generalization capability, capturing complex nonlinear relationships in the dataset more effectively. In particular, while traditional machine learning models rely on manually engineered features and predefined structures, deep learning approaches leverage hierarchical feature extraction, allowing for improved representation learning. The significantly higher predictive accuracy of the DNN model underscores its effectiveness in voltage prediction tasks for battery materials, particularly when dealing with high-dimensional feature spaces.
Figure 5: Performance comparison of traditional machine learning models SVR, RFR, and GBR for predicting the average voltage of battery materials.
Figure 6 present the results of a 10-fold cross-validation analysis, evaluating both the MSE and the R2 score for the training and testing sets. The low training MSE and the high training R2 (close to 1.0) indicate that the model effectively captures the underlying data patterns, while the consistently high testing R2 ($\sim$ 0.88) and stable testing MSE confirm its strong generalizability. To avoid overfitting, we employed feature selection techniques, regularization, and hyperparameter tuning to optimize the complexity of the model. In addition, the model was evaluated on an independent test set, ensuring its robustness against unseen data. These results demonstrate that the model achieves a balance between high predictive accuracy and generalization performance, making it well suited for battery voltage predictions.
Figure 6: Cross-validation performance of the DNN model: (Left) MSE for training and testing sets across 10 folds, with average values indicated by dashed lines. (Right) R² scores for training and testing sets, demonstrating model generalization.
Model validation is crucial for testing its excellent performance verification and evaluation. We validate our DNN model by considering experimentally observed and commercially available battery electrode material for both Li-ion and Na-ion batteries (see Table 2). In Figure 7, we show the predicted voltages against the DFT-calculated voltages for various electrode materials, demonstrating the accuracy and reliability of the model. The plot shows a strong correlation between the predicted and DFT-calculated voltages, with an R2 value of 0.991 and an MSE of 0.012V. The agreement between the predicted and experimental voltages is also excellent, with an average absolute deviation of 0.05V. This demonstrates the ability of the model to accurately predict voltages for various electrode materials, which is essential for the design of high-performance batteries.
@@
The list of materials used for the validation is shown in Table 2 along with voltage data obtained from DFT, our DNN model, and the experiment. The absolute deviation of the predicted voltage of our model with the experimental and DFT voltages is evaluated through the following expressions: $\Delta V_{1}=|V_{DFT}-V_{DNN}|$ and $\Delta V_{2}=|V_{expt.}-V_{DNN}|$, respectively, and is tabulated in Table 2.
@@
LiCoO2 4.169 3.808 3.803 0.005 0.297
LiFePO4 3.570 3.847 3.845 0.002 0.345
LiMnPO4 4.171 4.578 4.207 0.371 0.107
LiCoPO4 4.872 4.770 4.753 0.017 0.047
LiNiO2 3.8570 4.099 4.079 0.020 0.229
LiMn2O4 4.1573 4.577 4.545 0.032 0.395
LiTiS2 2.174 1.875 1.866 0.009 0.234
LiNiPO4 5.375 5.983 5.957 0.026 0.657
Li3V2(PO4)3 3.876 3.708 3.716 0.008 0.084
Na3V2(PO4)3 3.476 3.349 3.361 0.012 0.039
NaCoO2 2.877, 78 3.056 3.043 0.013 0.243
NaNiO2 3.079 3.315 3.277 0.038 0.277
NaTiO2 1.580 1.106 1.104 0.002 0.396
NaFePO4 3.081 3.072 3.046 0.026 0.046
@@
Our model is developed based on DFT data from battery materials and its associated features, ensuring that the deviation ($\Delta V_{1}$) between the predicted voltage of our DNN model and the calculated voltage by DFT remains minimal for nearly all validation materials. In contrast, a larger deviation ($\Delta V_{2}$) is observed when comparing the predicted values with the experimental data. This discrepancy is expected because the voltage of battery materials is influenced by various experimental conditions, such as humidity, temperature, and pressure, factors that are not considered in the DFT data set. Despite these challenges, our model demonstrates the ability to predict experimental voltages with minimal error, as evidenced by several cases presented in Table 2.
4.2 Feature importance
Figure 8 represents the analysis of the importance of characteristics for voltage prediction in battery materials that reveals that the most influential descriptors are related to the thermodynamic stability, electronic structure and redox behavior of compounds. The top two features, stability_discharge (MeV/atom) and stability_charge (MeV/atom), indicate the energetic stability of the material in both charge and discharge states. These features are crucial because they reflect the material’s ability to reversibly intercalate and deintercalate ions without significant structural degradation, directly affecting the voltage profile 82. The average p-valence electrons and the maximum number of unfilled p orbitals (max:num_p_unfilled) provide insight into the electronic structure, particularly the capacity of the material to accommodate electron transfer during redox reactions83. The minimum energy of the Ghosh electronegativity scale 84, 85 (min: en_ghosh) captures the material’s ability to stabilize anion-cation interactions, which is critical for determining the redox potential86. Furthermore, the standard deviation of the oxidation states reflects the variability in the oxidation behavior between the constituent elements, influencing the charge transfer dynamics31. Features such as fraction of s valence electrons, maximum ground-state magnetic moment (max:gs_mag_moment), and 5-norm30 (a measure of geometric or electronic structure variability) further contribute by characterizing bonding nature, magnetic properties, and structural stability, respectively. Lastly, the average van der Waals radius (ave:vdw_radius_uff) highlights the role of atomic size in determining ion mobility within the material87. These features collectively define the electrochemical and voltage characteristics of battery materials, which aids in the rational design of high-performance electrodes.
Figure 8: Key features affecting electrochemical performance.
4.3 ML-guided cathode design
@@
The design of new Na-ion battery cathode materials was carried out using template crystal structures of transition-metal layered oxides with the general formula NaMO2, where M represents transition or post-transition metals. These layered oxides serve as a fundamental structural framework as a result of their well-established electrochemical properties and stability in Na-ion battery applications. In this study, we focus on two widely studied structural types of layered transition-metal oxides. O3 and P2-type materials (see Figure 9). The O3-type materials exhibit trigonal symmetry with the R3mR-3m space group, while the P2-type materials possess hexagonal symmetry with the $P6_{3}/mmc$ space group. These two structural configurations were carefully considered in the design of new compositions due to their distinct sodium diffusion pathways and electrochemical characteristics.
@@
To generate novel compositions, we explored a wide range of 3d transition metals (Cr, Ti, V, Mn, Fe, Ni, Co, Cu, Zn) along with other selected metal species (Al, Mg, Sr), aiming to achieve an optimal balance of critical electrochemical properties. The primary selection criteria for these elements were phase stability, high voltage, long cycle life, efficient Na ion diffusion, and fast charging capability. The Supercell package 65 was utilized to systematically explore and enumerate various combinations of elements within the NaMO2 framework. This computational approach enabled the creation of an extensive library of potential compositions by generating numerous atomic substitutions within the host structure. The resulting data set includes a large number of unique chemical compositions for O3 and P2-type cathodes, each exhibiting distinct electrochemical and structural properties.
A comprehensive list of newly designed O3 and P2-type cathode materials is presented in Figure 10. The voltage predictions were obtained using our PyTorch-based DNN model, which has demonstrated superior accuracy in predicting the electrochemical performance of Na-ion cathodes. To further validate the stability and feasibility of these newly designed compositions, we performed first-principles DFT calculations. Tables 3 and 4 represent the newly designed material compositions with formation energy, assessing their thermodynamic stability and the likelihood of successful synthesis. A detailed discussion of the DFT results, including computed formation energies and structural analyses, is provided in the Supporting Information (SI) section.
@@
As summarized in Tables 3 and 4, all investigated O3 polymorphs exhibit formation energies that are 0.3–1.0 eV f.u.-1 lower (i.e., more negative) than those of their isochemical counterparts P2, indicating superior thermodynamic stability. This difference can be attributed to the distinct transition-metal–oxygen (TM–O) connectivity inherent to the two stacking sequences. In the O3 structure, each transition metal cation resides within a corner-sharing octahedron that also engages in extensive edge- and face-sharing with adjacent units. This results in a highly interconnected and six-fold coordinated TM-O framework. The enhanced orbital overlap within this dense network strengthens the overall TM–O bonding, making the creation of a sodium vacancy energetically more costly. In contrast, the P2 phases consist of prismatic TM–O units connected primarily through edge sharing, forming a more open and less interconnected framework. Consequently, the average TM–O interactions are weaker, leading to less negative formation energies.
4.4 Validation and analysis of ML-predicted materials
4.4.1 Voltage comparison (ML vs DFT)
Using DFT calculations within the GGA+U framework, we have validated the voltage predictions generated by our DNN model for specific compositions: O3-Na0.67Mn0.61Ni0.28Cr0.11O2 and P2-Na0.67Mn0.5Fe0.28Cu0.22O2. These compositions, highlighted in red in Figure 10(a, b), were selected as representative validation points to avoid the computational expense of performing DFT calculations for all possible compositions. The DNN predicted voltages for O3-Na0.67Mn0.61Ni0.28Cr0.11O2 and P2-Na0.67Mn0.5Fe0.28Cu0.22O2 are 3.79 and 3.76V, respectively, while the corresponding DFT-calculated voltages are 3.84 and 3.68V (see SI), as shown in Figure 10(a, b). The minimal errors of only 1.3% and 2.2% demonstrate excellent agreement between the two methods. This strong correlation highlights the robustness and reliability of our machine learning model in accurately predicting voltages while significantly reducing the computational cost associated with traditional first-principles calculations. The consistency between the DFT and DNN results underscores the potential of our model as a powerful tool for screening and predicting the electrochemical properties of novel sodium-ion battery materials.